OpenStudy (anonymous):

Write the equation of the line and draw the graph of the line through each pair of points given. Use either the slope-intercept equation or by using x- and y-intercepts >>4,6 and -1,2

OpenStudy (ganpat):

 Slope = (y2 - y1)/ (x2 -x1) = -4/-3= 4/3